OpenStudy (anonymous):

Amistre64 I don't get it! How do you get x=1 from that? I get x-1=x^2 and get stuck there. Help!

OpenStudy (amistre64):

-1 1 --- + --- = 0 ; how do we add fractions? we got to get them x*x x wearing the same pants otherwise they wont sit nice together right?

OpenStudy (amistre64):

we want to re dress that left one to match; we do that by multiplying it by a useful form of 1; lets use x/x

OpenStudy (amistre64):

-1 1(x) --- + ---- = 0 now we can add them x*x x(x)

OpenStudy (amistre64):

-1 + x ------ = 0 ; well, a fraction only = 0 when the top part is 0 x*x

OpenStudy (amistre64):

sooo: -1+x =0 ; +1 to each side x = 1

OpenStudy (anonymous):

oh my gosh! ok, i get that part where you said that a fraction can only = 0 when the numerator = 0. DUH, dumb me, I was trying to solve for x and you can't in the traditional way, can you?

OpenStudy (amistre64):

lol..... im gonna go out on a limb and say that that was the traditional way :)

OpenStudy (anonymous):

you kill me! Thank you again...now, back to calculus! If I can't get through the algebra in this, I am in big trouble, huh? ; )

OpenStudy (amistre64):

if you want to add an extra 'cross multiply' thats fine; then: -1+x = 0(x^2) -1 +x = 0

OpenStudy (anonymous):

Thanks again and hope you will be around for awhile! Maybe not just tonight but for the remainder of my calculus courses which is for the next 12 weeks!

OpenStudy (amistre64):

:) if not me; then there are smarter than mes all over the place ;)

OpenStudy (anonymous):

Maybe but you speak my language and not the language of mathspeak. THAT I don't get. I come here looking first for you, and you always seem to be here, so I'm sure I'll talk to you soon! Have a great night in Florida!

OpenStudy (amistre64):

thnx :) U2